Develop a formula to find the missing parts of the following triangle if the given parts are:
1. Angle B; side c
2. Angle A; side b
3. Angle B;Angle A
1
Expert's answer
2020-12-02T01:53:54-0500

1.Given "\angle B", side "c"

Find: "a-? b-? \angle A-?"

Solution:

Find the side "a":

"cos \angle B=\frac{a}{c}"

"a=c⋅cos\angle B"

Find the side "b":

"a^2+b^2=c^2"

"c^2⋅cos^2\angle B +b^2=c^2"

"b^2=c^2-c^2⋅cos^2\angle B"

"b^2=c^2⋅(1-cos^2\angle B)"

"b=c\sqrt{\smash[b]{1-cos^2\angle B}}"

Find "\angle A":

"\angle A=180\degree-\angle C-\angle B=180\degree-90\degree-\angle B="

"=90\degree-\angle B"

2.Given "\angle A", side "b"

Find: "a-?""c-?""\angle B-?"

Solution:

Find the side "c":

"cos \angle A=\frac{b}{c}"

"c⋅cos\angle A=b"

"c=\frac{b}{cos\angle A}"

Find the side "a":

"a^2+b^2=c^2"

"a^2+b^2=\frac{b^2}{cos^2\angle A}"

"a^2=\frac{b^2}{cos^2\angle A}-b^2"

"a^2=\frac{b^2-b^2⋅cos^2\angle A}{cos^2\angle A}"

"a^2=\frac{b^2}{cos^2\angle A}⋅(1-cos^2\angle A)"

"a=\frac{b}{cos\angle A}⋅\sqrt{\smash[b]{1-cos^2\angle A}}"

Find "\angle B":

"\angle B=180\degree-\angle A-\angle C=180\angle-\angle A- 90\degree="

"=90\degree-\angle A"

3.Given "\angle A, \angle B"

Find: "a-?" "b-?" "c-?"

Solution:

Let side "c=x", then find the side "b" :

"cos \angle A=\frac{b}{c}"

"b=c⋅cos\angle A"

"b=x⋅cos\angle A"

Find the side "a" :

"a^2+b^2=c^2"

"a^2+x^2⋅cos^2\angle A=x^2"

"a^2=x^2-x^2⋅cos^2\angle A"

"a^2=x^2(1-cos^2\angle A)"

"a=x⋅\sqrt{\smash[b]{1-cos^2\angle A}}"